Circular No. 5763

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 J. A. Phillips and S.R. Kulkarni, California Institute of Technology,

report: “Using the four-element millimeter array of the Owens Valley Radio Observatory, we have detected a compact, 98-GHz radio source at α = 9^{h}51^{m}19^{s}.1 ± 0^{s}.5, δ = +69^{o}15′27′′ ± 1′′ (equinox 1950.0), approximately coincident with the position of SN 1993J (IAUC 5731). Synthesis images (97.98 GHz) of a field centered at end figures 19^{s}.0, 26′′.5 were obtained on Apr. 11 and 14. On Apr. 11.24-11.52 UT, the flux density was 20 ± 5 mJy at end figures 18^{s}.8, 26′′; on Apr. 14.23-14.53, 29 ± 6 mJy at 19^{s}.3, 27′′. The synthesized beam was 5′′ × 4′′ on both dates and the source was unresolved. This marks the first detection of a supernova explosion at such high radio frequencies, which we intend to monitor at 1- to 2-week intervals.”

L. Morbidelli, TIRGO Observatory; and M. Kidger, Instituto de Astrofisica de Canarias, communicate: “We obtained near-infrared photometry of SN 1993J with the 1.5-m TIRGO Telescope in Gornergrat: Apr. 14.949 UT, J = 9.99 ± 0.07, H = 9.92 ± 0.08, K = 9.71 ± 0.13; 15.864, 10.12 ± 0.07, –, –; 15.937, 10.26 ± 0.10, 10.07 ± 0.04, 9.83 ± 0.03; 15.964, 10.13 ± 0.07, 10.04 ± 0.05, 9.78 ± 0.05. The data from Apr. 14.949 were taken in poor conditions. The data show that the infrared brightness has increased by 0.6-0.8 mag in all bands since Apr. 3.39 (IAUC 5757). The color changes are consistent with the development of strong hydrogen lines.”

GRB 930309 K. Hurley, Space Sciences Laboratory, University of California, Berke-

ley, on behalf of the Third Interplanetary Network team for triangulation of cosmic γ-ray bursts (M. Sommer, M. Boer, and M. Niel for Ulysses; G. Fishman, C. Kouveliotou, C. Meegan, and M. Brock for BATSE; T. Cline for Ulysses and BATSE), communicates: “An annulus of possible arrival directions has been derived for GRB 930309 (IAUC 5749, 5750, 5755) by timing the arrival of the event at the Ulysses and Compton-GRO (BATSE) spacecraft, which were separated by 1994 light-seconds. The annulus is centered at α = 22^{h}10^{m}24^{s}, δ = +14^{o}38′.7 (equinox 2000.0), with radius 36^{o}36′ ± 5′. The CGRO BATSE position is consistent with this narrow arc, but the radio source reported on IAUC 5750 and the infrared source reported on IAUC 5755 are at least 4^{o} from the timing arc and therefore are not related to the γ-ray burster.”

1993 April 16 Daniel W. E. Green